OpenStudy (anonymous):

3x-4y=-9 -6x+4y=-6 solve with elimination

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\\\huge\cal\color{red}W\color{orange}E\color{goldenrod}L\color{yellow}C\color{lightgreen}O\color{darkgreen}M\color{turquoise}E~~\color{royalblue}T\color{purple}O~~\color{#00bfff}{Open}\color{#11c520}{Study}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\)

OpenStudy (homeworksucks):

So let's write this out first to make our lives easier\[3x-4y=9\] \[-6x+4y=6\]

OpenStudy (homeworksucks):

Right away you can see that if we add the two equations to find one master equation, the y's cancel out

OpenStudy (homeworksucks):

So adding the two equations you get \[-3x+0y=15\] or simply\[-3x=15\]

OpenStudy (homeworksucks):

Now it's extremely easy to find x, just divide both sides by -3 to get -5

OpenStudy (homeworksucks):

What you want to do to find the y coordinate is to plug in your x value into one of your equations and solve for y

OpenStudy (homeworksucks):

Whoops the first equation is -9 not 9 haha my bad your x value is actually -3 when you change your values accordingly you get x=-3

OpenStudy (homeworksucks):

or rather just 5

OpenStudy (homeworksucks):

Anyway so once you plug in your x value of 5 into let's say the first equation (it can be either of the two) you get \[3(5)-4y=-9\]

OpenStudy (homeworksucks):

subtract 15 from both sides to get -4y=-24

OpenStudy (homeworksucks):

then solve for y, which is 6

OpenStudy (homeworksucks):

So your solution is (x,y)=(5,6)